OpenStudy (anonymous):

Eliminate the parameter. x = t + 4, y = t^2

myininaya (myininaya):

solve x=t+4 for t then replace the t in y=t^2 and then done

myininaya (myininaya):

@notmathnerd do you understand?

OpenStudy (anonymous):

yes I do thank you

OpenStudy (anonymous):

would it be y=x^2 - 16 ? @myininaya

myininaya (myininaya):

hmmm not correct but i think it is because you forgot a simple algebraic rule (and don't worry a lot of people forget it)

myininaya (myininaya):

(a-b)^2 doesn't meant a^2-b^2

myininaya (myininaya):

(a-b)^2 means (a-b)(a-b)

myininaya (myininaya):

so what am i saying is i'm pretty sure you got this far y=(x-4)^2 but you didn't expand the (x-4)^2 correctly that is all

OpenStudy (anonymous):

So the answer is actually x^2 - 8x + 16?

myininaya (myininaya):

the answer is actually y=x^2-8x+16

OpenStudy (anonymous):

ok thank you